On January 1, Year 1, Bryson Company obtained a $39,000, four-year, 10% installment note from Campbell Bank. The note requires annual payments of $12,303, beginning on December 31, Year 1.

a. Prepare an amortization table for this installment note, similar to the one presented in Exhibit 4.

Note: Round the computation of the interest expense to the nearest whole dollar. Enter all amounts as positive numbers. In Year 4, round the amount in the Decrease in Notes Payable column either up or down to ensure that the Carrying Amount zeroes out.

Amortization of Installment Notes
Year Ending
December 31

January 1
Carrying Amount

Note Payment
(Cash Paid)
Interest Expense
(10% of January 1
Note Carrying
Amount)

Decrease in
Notes Payable

December 31
Carrying Amount
Year 1 $ $ $ $ $
Year 2
Year 3
Year 4 0
$ $ $

b. Journalize the entries for the issuance of the note and the four annual note payments.

Note: For a compound transaction, if an amount box does not require an entry, leave it blank. For the Year 4 entry (due to rounding), adjust Notes Payable up or down to ensure that debits equal credits.

Year 1 Jan. 1
Year 1 Dec. 31
Year 2 Dec. 31
Year 3 Dec. 31
Year 4 Dec. 31

c. How will the annual note payment be reported in the Year 1 income statement?
  of $ would be reported on the income statement.
Answer #1

Answers

  • [A]

Year Ending Dec 31

Jan 1 Carrying amount

Notes Payament Cash paid

Interest Expense

Decrease in Notes Payable

Dec 31 Carrying Amount

Year 1

$39,000

$12,303

$3,900

$8,403

$30,597

Year 2

$30,597

$12,303

$3,060

$9,243

$21,354

Year 3

$21,354

$12,303

$2,135

$10,168

$11,186

Year 4

$11,186

$12,303

$1,117

$11,186

$0

--Working

Year Ending Dec 31

Jan 1 Carrying amount

Notes Payament Cash paid

Interest Expense

Decrease in Notes Payable

Dec 31 Carrying Amount

Year 1

39000

12303

=39000*10%

=12303-3900

=39000-8403

Year 2

=39000-8403

12303

=30597*10%

=12303-3060

=30597-9243

Year 3

=30597-9243

12303

=21354*10%

=12303-2135

=21354-10168

Year 4

=21354-10168

12303

=12303-11186

11186

=11186-11186

  • [B]

Date

Accounts title

Debit

Credit

Year 1 Jan 1

Cash

$39,000

Notes Payable

$39,000

Year 1 Dec 31

Notes Payable

$8,403

Interest Expense

$3,900

Cash

$12,303

Year 2 Dec 31

Notes Payable

$9,243

Interest Expense

$3,060

Cash

$12,303

Year 3 Dec 31

Notes Payable

$10,168

Interest Expense

$2,135

Cash

$12,303

Year 4 Dec 31

Notes Payable

$11,186

Interest Expense

$1,117

Cash

$12,303

  • [C]

$ 3,900 of $ 12303 will be reported on Income Statement as Interest Expense.